What is the equations in slope intercept form for the line that contains (1,1) and (3,5)
Answer by lenny460(1073)

What is the equations in slope intercept form for the line that contains (1,1) and (3,5)

First let us find the slope of the straight line.
(y2 - y1) / (x2 - x1)
= (5 - 1) / (3 - 1)
5 - 1 = 4
3 - 1 = 2
4/2 = 2
The Slope (m) = 2

Now let us fine the slope intercept form:
y - y1 = m(x - x1)
y - 5 = 2(x - 3)
2 * x = 2x
2 * - 3 = -6
Therefore:
y - 5 = 2x - 6
Add 5 to each side of the equation.
y - 5 + 5 = 2x - 6 + 5
Combine like terms:
-5 + 5 = 0
-6 + 5 = -1
Therefore:
y = 2x - 1
Slope = 2
y-intercept = -1

Therefore the slope intercept form
of this line is:
y = 2x - 1



Lennox Obuong
Algebra Student
Email: obuong3@aol.com